OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form? y = negative one-fourthx + 2 y = one-fourthx + 2 y = 4x + 2 y = –4x + 3

OpenStudy (anonymous):

do you any clues on how to solve it?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok so the first thing we need to notice is that they give us one point (0,2) and its perpendicular to the line of y = 4x +3. so to find the equation of the other line we will have to use the point given (0,2) and also find something else to help us get to the final answer, and it is the gradient/slope.

OpenStudy (anonymous):

They tell us its perpendicular, this means that m1 x m2 = -1 m1=gradient one m2=gradient two from the equation y = 4x + 3 what is the gradient equal to?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

ok, you have to remember this.

OpenStudy (anonymous):

For the equation y = 4x + 3 and any other equation rearranged in the form y = ax + c the gradient/slope represented by the symbol m is equal to whatever is in front of the x, in this case it is 4.

OpenStudy (anonymous):

y = 4x + 3 m(gradient/slope)= 4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes. i understand now

OpenStudy (anonymous):

ok, good.

OpenStudy (anonymous):

lets try one quick example first, so what would the gradient of y=10x + 2 be?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

yes, that's great.

OpenStudy (anonymous):

so the answer is c? right?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

a*?

OpenStudy (anonymous):

ok, I am not going to tell you the answer yet as guessing is not the right way to understanding.

OpenStudy (anonymous):

so we have one point given (0,2) and the unknown line has to be perpendicular to the line of y=4x + 3. m1=gradient of first line (this is known from y=4x + 3, which is 4) m2=gradient of second line (unknown) m1 x m2 = -1 4 x m2 = -1 m2 = -1/4 y=mx + c y= -1/4x + c (0,2) 2=-1/4(0) + c 2= 0 + c c = 2 therefore the answer is y=-1/4x + 2 which is answer A.

OpenStudy (anonymous):

it all makes sense now. ill make sure to remember this next time. Thanks @JayDS

OpenStudy (anonymous):

your welcome, I just want to make sure you really understand it.

OpenStudy (anonymous):

i appriciate that. i understand now. thanks again!